Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 366     Entry time: 24 Sep 2023, 07:49
Author: Natalia, Justina, Giacomo, Sergey, Ben, Eugenio, Teguh 
Group: SCS 
Subject: summary of 23.09 

Around 6:55 there was a power glitch. Power in the experimental hutch and rack room went down in SASE3, error message in Xray interlock system and no power for PP laser.

During morning TS turned the power on and interlock issues were solved by DESY. SCS could restart large part of SCS systems, however got help of DOC only after 2:30p.m. Similarily, PP laser was needed help of DOC and got this help only around 2p.m. The afternoon shift was mainly bringing the instrument in operation with helps of DOC. Around 9p.m. we realized that PP laser is not in operational state because of large fluctuations, this was not communicated to SCS at all. After calling the Laser group, SCS was told that the laser group will work on recover of the laser on Sunday morning. Without laser we could not continue user program, therefore  at least 24h of user operation were lost.

 

ELOG V3.1.4-7c3fd00